# Depth-Structured Music Recurrence: Budgeted Recurrent Attention for Full-Piece Symbolic Music Modeling

### v1

@@ III-C Full-attention approximation of a vanilla Transformer as a reference model
$(K_{<t}^{\star},V_{<t}^{\star})=\bigl(K_{<t}(\theta_{t}),\ V_{<t}(\theta_{t})\bigr).$ (3)
@@
$(K_{<t},V_{<t})=(K_{<t}^{\star},V_{<t}^{\star})+(\Delta K_{<t},\Delta V_{<t}),$ (4)
where $(\Delta K_{<t},\Delta V_{<t})$ captures the discrepancy due to reusing cached states produced by earlier parameters.
On the MAESTRO piano performance dataset, with a maximum context length $L_{\max}$ and segment length $L_{\mathrm{seg}}$, each piece is processed for at most $T\leq\lceil L_{\max}/L_{\mathrm{seg}}\rceil$ recurrent steps, so the cached history concatenates states produced by no more than $T$ previous steps. In our experiments, the best early-stopping checkpoint typically occurs after roughly $3$ million training steps; since $T$ is orders of magnitude smaller (in our setting, $T\leq 32$), the step-to-step parameter mismatch over the cached history appears small and does not impede convergence.
Removing time-axis truncation provides a practical approximation to full self-attention in a vanilla Transformer, while retaining the computational benefits of segment-wise processing; Appendix A-B analyzes this approximation.
III-D Depth-Structured Music Recurrence (DSMR) with layer-wise horizon
@@
III-E Two-scale DSMR: long horizons in lower layers + a uniform short horizon elsewhere
@@
III-F Gate-guided layer selection as a comparison variant
In binary-horizon schedules, besides heuristic choices of $\mathcal{L}_{\mathrm{mem}}$, we also evaluate a gate-guided variant that uses learnable gates to provide a data-driven signal for memory-layer selection. Concretely, we perform a gate-induced intervention by training a gated model, ranking layers by the mean gate usage $\bar{g}^{(\ell)}$ to form the top-$B$ set $\mathcal{L}_{\mathrm{mem}}$, and retraining the resulting selective-retention model; we then compare against sliding-window, uniform, and random choices of $\mathcal{L}_{\mathrm{mem}}$ under the same $B$.
[…]
We also evaluate a late-stage discretization variant with the Straight-Through Estimator (STE) [4, 10]: after 250250k steps, each gate is binarized to $\{1.0,\,0.0\}$ (open/closed), and the resulting binary pattern defines an alternative $\mathcal{L}_{\mathrm{mem}}$ for comparison under the same $B$.
IV Experimental Setup
@@
IV-A Datasets and tokenization
We conduct experiments on the MAESTRO dataset, a large-scale collection of virtuosic piano performances with MIDI note events [20]. We filter the dataset to retain only samples whose tokenized lengths fall between 1024 and 32768 tokens. We use the train/validation/test split produced by our preprocessing pipeline (0.89/0.10/0.01), and train an autoregressive model over tokenized symbolic event sequences derived from MIDI.
@@ IV-A Datasets and tokenization
IV-B Model and recurrence configuration
We modify a Transformer-XL style recurrent Transformer as backbone (Sec. A-A). Long performances are treated as continuous streams and trained recurrently with sliding windows (segments) of length $s=1024$, while carrying a stop-gradient key–value (KV) history cache across segments. This enables each segment to attend to historical context beyond its local span under limited compute.
The model has $L=18$ layers, hidden size $d=1024$, $h=16$ attention heads (with $d_{\mathrm{head}}=64$), and FFN size $d_{\mathrm{ff}}=4d=4096$. We implement Depth-Structured Music Recurrence (DSMR)’s layer-wise temporal receptive fields (Sec. III) by specifying per-layer memory horizons $\{m^{(\ell)}\}_{\ell=1}^{L}$ for the recurrent KV cache. Each layer $\ell$ maintains a recurrent KV memory capped at $m^{(\ell)}$ tokens, with a per-layer maximum of $m_{0}=31744$ tokens. We fix the overall recurrent-state budget to $M_{\mathrm{tot}}=95232$ (i.e., $31744\times 3$), so varying $\{m^{(\ell)}\}$ redistributes where along depth long-range history is accessible under the same total memory.
@@ IV-B Model and recurrence configuration
We use the two-scale DSMR schedule in Eq. (7). We set $L_{\mathrm{long}}=1$ and $L_{\mathrm{short}}=17$ (so $L=18$), where the bottom $L_{\mathrm{long}}$ layer uses a long horizon $m_{\mathrm{long}}=31744$, and the remaining $L_{\mathrm{short}}$ layers share a uniform short horizon $m_{\mathrm{short}}=3735$.
IV-C Other model variants
We study depth-wise allocation of recurrent memory by instantiating DSMR with several horizon schedules under matched budgets. All variants share the same backbone architecture and training setup; they differ only in the per-layer horizon vector $\{m^{(\ell)}\}_{\ell=1}^{L}$ and (optionally) a lightweight memory gate. These variants include selective retention (binary on/off horizons), soft gating, progressive depth-dependent caps, a reversed two-scale schedule, and a Perceiver-AR-like schedule; their memory-horizon schedules (layer-wise horizon allocations) are illustrated in Fig. 3. We report their performance alongside the final two-scale DSMR model for direct comparison.
@@ IV-C Other model variants
As a comparison variant in the binary-horizon setting, we attach a soft gate $g^{(\ell)}_{t}\in[0,1]$ to scale only the historical KV memory in each retained layer (Eq. (9)). We train the gated model, rank layers by mean gate usage $\bar{g}^{(\ell)}$ to select the top-$B$ layers as $\mathcal{L}_{\mathrm{mem}}$, and retrain the resulting selective-retention model for comparison against sliding-window, uniform, and random $\mathcal{L}_{\mathrm{mem}}$ under the same $B$. We also evaluate an STE discretization variant [4, 10] by binarizing gates after 250250k steps to form an alternative $\mathcal{L}_{\mathrm{mem}}$.
@@
Perceiver AR achieves long-context modeling by performing cross-attention between the current segment queries and a long key–value memory: the queries come from the short attention segment, while the K/V states are computed by projecting token embeddings from the concatenated context sequence. Appendix A-C provides a detailed explanation. Motivated by this design, our Perceiver-AR-like schedule retains full-length (maximum-horizon) recurrent memory only at the shallowest layer, and sets the recurrent horizon of all higher layers to zero (i.e., no cross-segment memory), so higher layers perform only within-segment self-attention.
IV-D Training details
@@
with warmup $w=10000$ and $\mathrm{LR}_{0}=1.0$. This gives a peak learning rate $\mathrm{lr}_{\max}=\mathrm{LR}_{0}/\sqrt{d\,w}=3.125\times 10^{-4}$ for $d=1024$. We take one optimizer step per sliding window. We do not apply gradient clipping and use a cross-attention prefix dropout of $0.1$. All experiments run on CUDA-enabled consumer-grade GPUs. We repeat each key configuration with random seeds and report the mean and standard deviation.
IV-E Evaluation protocol and metrics
@@
Unless stated otherwise, we report Best Val PPL at the checkpoint with the lowest validation loss. We also report the compute efficiency (throughput) and cost (wall-clock time and peak memory) in absolute terms for both our models and the full-attention Transformer-XL reference, to characterize the budget–quality tradeoff.
TABLE I: Key hyperparameters used in our experiments.
@@
Seeds NN 8
V Results
We organize results around our central claim that Depth-Structured Music Recurrence (DSMR) achieves full-piece long-context music modeling through segment-level recurrence with detached cross-segment states, budgeted across depth by a layer-wise memory-horizon schedule. Accordingly, we present three subsections. First, we benchmark DSMR against representative long-context baselines and recurrent strategies under matched recurrent-state budgets, reporting both quality (Best Val PPL) and efficiency (tokens/s, peak GPU memory, and time to the best validation checkpoint). Second, we analyze binary-horizon DSMR (selective retention) to test whether performance depends on which layers retain memory or primarily on the memory-depth budget, finding substantial layer substitutability under recurrence. Third, we study learned memory-usage gates as a candidate signal for choosing memory-carrying layers, showing that while gate preferences are non-uniform, gate-guided interventions do not consistently improve over budget-matched DSMR schedules.
V-A Main comparison and efficiency under a matched recurrent-state budget
We compare recurrent strategies and report both quality and training efficiency. Let $m_{0}=31744$ denote the maximum per-layer horizon (Sec. IV-B). As an unconstrained reference, we report: (i) a full-memory baseline that retains essentially all cross-segment history (i.e., uses the maximum horizon at every layer). For matched-budget comparisons indexed by $B=3$, we constrain the total horizon as $\sum_{\ell=1}^{L}m^{(\ell)}=M_{\mathrm{tot}}=B\,m_{0}$. Under this constraint, we compare: (ii) binary-horizon DSMR (selective retention) that assigns $m^{(\ell)}\in\{0,m_{0}\}$ with exactly $B$ active layers, and (iii) DSMR redistribution schedules that allocate memory across all depths via $\{m^{(\ell)}\}$ under the same total budget. Our final model is the two-scale DSMR schedule (long horizons in lower layers, uniform short horizons elsewhere).
For quality, we report Best Val PPL, evaluated at the checkpoint that achieves the lowest validation loss (early stopping by validation). Figure 4 summarizes the quality–budget tradeoff. Under comparable memory usage ($\sim$6–8,GB), the two-scale DSMR schedule achieves the lowest Best Val PPL among non-full-attention methods (5.96, evaluated at the best validation checkpoint), outperforming gate-guided selective retention (6.53) and the Perceiver-AR-like reference (6.54), while the full-attention Transformer-XL reference attains 5.98 at substantially higher memory cost (15.5,GB). This supports the key takeaway: under a fixed recurrent-state budget, enforcing layer-wise temporal receptive fields (i.e., allocating horizons across depth) is more effective than searching over a small subset of memory-carrying layers.
@@ V-A Main comparison and efficiency under a matched recurrent-state budget
A key motivation is to enable long-context symbolic music modeling under constrained compute, so in addition to the peak GPU memory reported in Figure 4, we additionally report training efficiency: tokens processed per second, and wall-clock time to the best validation checkpoint. Table II summarizes efficiency metrics. Compared with the full-attention reference, DSMR reduces the KV source length in a subset of attention layers by enforcing layer-wise horizons, thereby controlling the recurrent-state footprint. Using over 50% less GPU memory than the full-attention reference, the two-scale DSMR schedule achieves a strong quality–efficiency tradeoff: it stays within the fast-training regime of selective methods, and matches (slightly surpassing) the full-attention Best Val PPL.
TABLE II: Efficiency metrics on MAESTRO. All runs see 15,098,82515{,}098{,}825 training tokens. Percentages in parentheses are relative to the full-attention reference. Bold indicates the best value among non-full-attention methods.
@@
Full-attention Transformer-XL reference 7,618 15.5 6.06 5.98
@@
Gate-guided selective retention DSM 11,141 (+46.2%) 7.8 (-49.7%) 4.14 (-31.7%) 6.53 (+9.2%)
Progressive Ascending DSM 10,540 (+38.4%) 7.0 (-54.5%) 4.38 (-27.7%) 6.15 (+2.9%)
Progressive Descending DSM 10,556 (+38.6%) 7.1 (-54.3%) 4.37 (-27.9%) 6.33 (+5.9%)
Reverse Two-scale DSM 10,324 (+35.5%) 6.3 (-59.1%) 4.47 (-26.2%) 6.01 (+0.5%)
@@
V-B Model variant: selective retention (binary-horizon DSMR)
We next evaluate the binary-horizon DSMR variant (selective retention) to test whether performance depends on which depths carry recurrent memory when the depth budget is fixed, while keeping the nonzero horizon at the maximum per-layer horizon in every retained layer. At a fixed depth budget $B$, we compare multiple retained-layer subsets $\mathcal{L}_{\mathrm{mem}}$ (sliding-window, uniform spacing, and random subsets), where $m^{(\ell)}=m_{0}$ for $\ell\in\mathcal{L}_{\mathrm{mem}}$ and $m^{(\ell)}=0$ otherwise.
Table II shows that, across the Selective retention DSMR settings, performance varies little under different subset patterns. These observations indicate strong structural substitutability in the binary-horizon regime: within our setting, performance is largely determined by how many layers have nonzero temporal receptive fields (the budget $B$), rather than which layers they are.
V-C Model variant: gating reveals preference but not transferable importance
We then study whether learned memory usage provides actionable guidance for allocating memory retention across layers. When we equip memory-carrying layers with a learnable gate $g^{(\ell)}$ (Sec. IV-C), gate dynamics exhibit overall shrinkage together with increasing differentiation across depth during training: $\mathrm{KL}_{\text{uniform}}$ increases over steps, while gate entropy and the effective number of active layers decrease (Figure 5).
@@ V-C Model variant: gating reveals preference but not transferable importance
These trends persist even without an explicit global gate budget or binarization regularization. In late training (after 250k steps), applying a Straight-Through Estimator (STE) [4, 10] drives gates toward a near-discrete regime (open/closed) (Figure 7), while without STE gates remain continuous but still separate across depth (Figure 6).
@@
Figure 8 reports post-training layer-wise mean gates $\bar{g}^{(\ell)}$. Although gate values are clearly non-uniform across depth, using gate ranking to choose $\mathcal{L}_{\mathrm{mem}}$ and retraining under the same depth budget does not yield consistent gains over simple heuristic layer selections (Table II). This suggests that gate magnitudes reflect coordination/credit assignment in the original computation graph, rather than transferable causal importance under structural intervention on the layer-wise temporal receptive fields.
@@
VI Discussion
This section contextualizes our findings from four complementary angles. First, we clarify why access to long, piece-level context remains essential for generation where global musical form matters. Second, we argue that when the full context is already available, deterministic “deep” summaries do not improve the Bayes-optimal log-loss beyond having that context available at a shallow stage. Third, we examine how detached recurrent states reused across segments can become stale, yet do not necessarily compromise optimization stability under tight memory budgets. Finally, we discuss practical implications for long-context symbolic music modeling under constrained resources and outline limitations and scope conditions under which our conclusions may weaken.
VI-A Longer context as a richer information structure
A natural concern is that longer context may introduce irrelevant tokens and worsen validation perplexity. However, in terms of the optimal achievable validation log-loss over a model class, additional context cannot hurt as long as the longer-context family strictly contains the shorter-context one (e.g., it can implement predictors that ignore the extra tokens).
Equivalently, if a model class that conditions on context length $L_{2}>L_{1}$ subsumes the predictors available under $L_{1}$, then the best achievable validation NLL (and thus validation perplexity, since $\mathrm{PPL}=\exp(\mathrm{NLL})$) under $L_{2}$ is no worse than under $L_{1}$.
This is an instance of the “more information cannot hurt” principle: a richer information structure weakly improves the best attainable risk under log-loss [5, 13, 17]. Crucially, the claim is about an infimum over the hypothesis class rather than any single non-convex training run; optimization and finite-data effects can still prevent realized validation PPL from improving. Nevertheless, for generation where global musical form matters, piece-level context remains essential.
In our DSMR experiments, we therefore use a receptive-field schedule that includes at least one layer with an effectively unbounded horizon, ensuring that every token can attend to all preceding tokens. This design aims to preserve piece-level information for downstream generation, so the model can learn complete musical structure, including motif repetition and developmental variation.
VI-B Deterministic deep processing cannot improve the Bayes-optimal log-loss
This section formalizes the limit of using “deeper” context representations. Let $X$ denote the available context (conditioning information) and $Y$ the prediction target (the next token/event). Let $Z=g(X)$ be a deterministically processed representation of $X$, without injecting extra information (e.g., retrieval). Then $Y\to X\to Z$ is a Markov chain, so by the data-processing inequality, $I(Y;Z)\leq I(Y;X)$, equivalently $H(Y\mid X)\leq H(Y\mid Z)$ [11].
Under log-loss, the Bayes-optimal expected negative log-likelihood equals conditional entropy, so the best achievable NLL using only $Z$ cannot be lower than that using the full context $X$. Equality is possible only when $Z$ is sufficient for predicting $Y$ (e.g., when $p(Y\mid X)=p(Y\mid Z)$, or when $g$ is invertible).
Therefore, improvements in realized validation PPL from operating on “deep” representations must come from factors other than surpassing the Bayes information limit–such as inductive bias, optimization/regularization, or additional injected information [17, 11].

### v2

@@ 3.3 Full-attention approximation of a vanilla Transformer as a reference model
$(K_{\theta},V_{\theta})=\bigl(K(\theta),\ V(\theta)\bigr).$ (3)
@@
$(K_{\text{cache}},V_{\text{cache}})=(K_{\theta},V_{\theta})+(\Delta K,\Delta V),$ (4)
where $(\Delta K,\Delta V)$ captures the discrepancy due to reusing cached states produced by earlier parameters. We analyze the implications of this staleness for optimization stability in Sec. 6.2.
Removing time-axis truncation provides a practical approximation to full self-attention in a vanilla Transformer, while retaining the computational benefits of segment-wise processing; Appendix A.2 analyzes this approximation.
3.4 Layer-wise horizon schedules
@@
3.5 Two-scale schedule
@@
4 Experimental Setup
@@
4.1 Datasets and tokenization
We conduct experiments on the MAESTRO dataset, a large-scale collection of virtuosic piano performances with MIDI note events Hawthorne et al. [2019]. We filter the dataset to retain only samples whose tokenized lengths fall between 1024 (the segment length) and 32768 tokens (the maximum context length), ensuring every piece spans at least one full segment and fits within the recurrence horizon. We use the train/validation/test split produced by our preprocessing pipeline (0.89/0.10/0.01), and train an autoregressive model over tokenized symbolic event sequences derived from MIDI.
@@ 4.1 Datasets and tokenization
4.2 Model and recurrence configuration
The model has $L=18$ layers, hidden size $d=1024$, $h=16$ attention heads (with $d_{\mathrm{head}}=64$), and FFN size $d_{\mathrm{ff}}=4d=4096$. Each piece is streamed left-to-right in segments of length $s=1024$ using the recurrent backbone described in Sec. 3.1, with per-layer memory horizons $\{m^{(\ell)}\}_{\ell=1}^{L}$ for the recurrent KV cache. Each layer $\ell$ maintains a recurrent KV memory capped at $m^{(\ell)}$ tokens, with a per-layer maximum of $m_{0}=31744$ tokens. We fix the overall recurrent-state budget to $M_{\mathrm{tot}}=95232$ (i.e., $31744\times 3$), so varying $\{m^{(\ell)}\}$ redistributes where along depth long-range history is accessible under the same total memory.
@@ 4.2 Model and recurrence configuration
We use the two-scale DSMR schedule in Eq. (7). We set $L_{\mathrm{long}}=1$ and $L_{\mathrm{short}}=17$ (so $L=18$), where the bottom $L_{\mathrm{long}}$ layer uses a long horizon $m_{\mathrm{long}}=31744$, and the remaining $L_{\mathrm{short}}$ layers share a uniform short horizon $m_{\mathrm{short}}=3734$.
4.3 Variant schedules
We study depth-wise allocation of recurrent memory by instantiating DSMR with several horizon schedules under matched budgets. All variants share the same backbone architecture and training setup; they differ only in the per-layer horizon vector $\{m^{(\ell)}\}_{\ell=1}^{L}$ and (optionally) a lightweight memory gate. Fig. 3 illustrates the layer-wise horizon allocations for all variants.
@@ 4.3 Variant schedules
[…]
@@
Perceiver AR achieves long-context modeling by performing cross-attention between the current segment queries and a long key–value memory: the queries come from the short attention segment, while the K/V states are computed by projecting token embeddings from the concatenated context sequence Hawthorne et al. [2022]. Appendix A.3 provides a detailed explanation. Motivated by this design, our Perceiver-AR-like schedule retains full-length (maximum-horizon) recurrent memory only at the shallowest layer, and sets the recurrent horizon of all higher layers to zero (i.e., no cross-segment memory), so higher layers perform only within-segment self-attention.
4.4 Training details
@@
with warmup $w=10000$ and $\mathrm{LR}_{0}=1.0$. This gives a peak learning rate $\mathrm{lr}_{\max}=\mathrm{LR}_{0}/\sqrt{d\,w}=3.125\times 10^{-4}$ for $d=1024$. We take one optimizer step per segment. We do not apply gradient clipping and use a cross-attention prefix dropout of $0.1$. All experiments run on CUDA-enabled consumer-grade GPUs.
4.5 Evaluation protocol and metrics
@@
Unless stated otherwise, we report Best Val PPL at the checkpoint with the lowest validation loss. We also report the compute efficiency (throughput) and cost (wall-clock time and peak memory) in absolute terms for both our models and the full-memory (untruncated-horizon) Transformer-XL-style recurrent reference, to characterize the budget–quality tradeoff.
Table 1: Key hyperparameters used in our experiments.
@@
5 Results
We present results in three parts: (i) quality and efficiency of DSMR schedules against baselines under matched recurrent-state budgets; (ii) layer substitutability under binary-horizon (selective retention) schedules; and (iii) whether learned memory-usage gates provide actionable guidance for layer selection.
5.1 Quality–efficiency tradeoff
We compare all DSMR schedules and baselines under the matched budget $M_{\mathrm{tot}}=B\,m_{0}$ (B=3B{=}3), with a full-memory reference as an unconstrained upper bound (Sec. 4.2). Figure 4 summarizes the quality–budget tradeoff. Under comparable memory usage ($\sim$6–8 GB), the two-scale DSMR schedule achieves the lowest Best Val PPL among budgeted methods (5.96), outperforming gate-guided selective retention (6.53) and the Perceiver-AR-like reference (6.54), while the full-memory reference attains 5.98 at substantially higher memory cost (15.5 GB). Under a fixed recurrent-state budget, redistributing horizons across all depths is more effective than concentrating memory in a small subset of layers.
@@ 5.1 Quality–efficiency tradeoff
Table 2 additionally reports training efficiency. The two-scale DSMR schedule uses 59.1% less GPU memory and trains 26.4% faster (wall-clock time to best checkpoint) than the full-memory reference, while matching its Best Val PPL (5.96 vs. 5.98).
Table 2: Efficiency metrics on MAESTRO. All runs see 15,098,82515{,}098{,}825 training tokens. Percentages in parentheses are relative to the full-attention reference. Bold indicates the best value among non-full-attention methods.
@@
Full-memory TXL-style reference 7,618 15.5 6.06 5.98
@@
Gate-guided selective retention DSMR 11,141 (+46.2%) 7.8 (-49.7%) 4.14 (-31.7%) 6.53 (+9.2%)
Progressive Ascending DSMR 10,540 (+38.4%) 7.0 (-54.5%) 4.38 (-27.7%) 6.15 (+2.9%)
Progressive Descending DSMR 10,556 (+38.6%) 7.1 (-54.3%) 4.37 (-27.9%) 6.33 (+5.9%)
Reverse Two-scale DSMR 10,324 (+35.5%) 6.3 (-59.1%) 4.47 (-26.2%) 6.01 (+0.5%)
@@
5.2 Layer substitutability under selective retention
We evaluate the binary-horizon DSMR variant (selective retention) to test whether performance depends on which depths carry recurrent memory when the depth budget $B$ is fixed. We compare sliding-window, uniform-spacing, and random retained-layer subsets $\mathcal{L}_{\mathrm{mem}}$, where $m^{(\ell)}=m_{0}$ for $\ell\in\mathcal{L}_{\mathrm{mem}}$ and $m^{(\ell)}=0$ otherwise.
Table 2 shows that performance varies little across subset patterns (Best Val PPL: sliding-window 6.51, uniform 6.52, random 6.53). This indicates strong structural substitutability: within our setting, performance is largely determined by the number of layers with nonzero horizons (the budget $B$), rather than which specific layers carry memory.
5.3 Gate preferences versus transferable importance
We then study whether learned memory usage provides actionable guidance for allocating memory retention across layers. When we equip memory-carrying layers with a learnable gate $g^{(\ell)}$ (Sec. 4.3), gate dynamics exhibit overall shrinkage together with increasing differentiation across depth during training: $\mathrm{KL}_{\text{uniform}}$ increases over steps, while gate entropy and the effective number of active layers decrease (Figure 5).
@@ 5.3 Gate preferences versus transferable importance
These trends persist even without an explicit global gate budget or binarization regularization. In late training (after 250k steps), applying a Straight-Through Estimator (STE) Bengio et al. [2013], Courbariaux et al. [2015] drives gates toward a near-discrete regime (open/closed) (Figure 7), while without STE gates remain continuous but still separate across depth (Figure 6).
@@
Figure 8 reports post-training layer-wise mean gates $\bar{g}^{(\ell)}$. Although gate values are clearly non-uniform across depth, using gate ranking to choose $\mathcal{L}_{\mathrm{mem}}$ and retraining under the same depth budget does not yield consistent gains over heuristic layer selections (Table 2). This suggests that gate magnitudes reflect coordination/credit assignment in the original computation graph, rather than transferable causal importance under structural intervention on the layer-wise temporal receptive fields.
@@
6 Discussion
We contextualize our findings from four angles: an information-theoretic perspective on context depth, state staleness under recurrence, practical implications, and limitations.
6.1 Information-theoretic perspective on context depth
Two information-theoretic observations underpin the DSMR design.
First, longer context cannot worsen the optimal achievable log-loss. If a model class conditioning on context length $L_{2}>L_{1}$ subsumes the predictors available under $L_{1}$, then the best achievable validation NLL under $L_{2}$ is no worse than under $L_{1}$—an instance of the “more information cannot hurt” principle Blackwell [1953], Dawid [1984], Gneiting and Raftery [2007]. This motivates retaining at least one layer with an effectively unbounded horizon, so every token can attend to all preceding tokens and piece-level structure is preserved.
Second, deterministic deep processing cannot improve the Bayes-optimal log-loss. Let $X$ denote the context, $Y$ the prediction target, and $Z=g(X)$ a deterministic representation of $X$. Then $Y\to X\to Z$ forms a Markov chain, so by the data-processing inequality $H(Y\mid X)\leq H(Y\mid Z)$ Cover and Thomas [2006]: the best achievable NLL using $Z$ alone cannot beat that using the full context $X$. Any realized improvement from deeper representations must therefore come from inductive bias or optimization effects, not from surpassing the information limit Gneiting and Raftery [2007], Cover and Thomas [2006].